OpenStudy (dmezzullo):

The table below shows the fees to check bags on a certain airline. A 2-column table with 'Type of Bag' and 'Fee per Bag' as the column headings. The first checked bag costs $20, the second checked bag costs $25, and extra bags cost $50 each If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all? Numerical Answers Expected!

OpenStudy (dmezzullo):

@AriPotta

OpenStudy (anonymous):

C = 20 + 25 + 50 ( x - 2) = 145 C = 45 + 50 (x - 2) = 145 50 (x - 2) = 100 x - 2 = 2 x = 4

OpenStudy (dmezzullo):

@satellite73

OpenStudy (dmezzullo):

TYVM